# Characterization of Josephson Junction Aging and Annealing Under Different Environments

### v1

@@ Abstract
Understanding the aging behavior of Josephson junctions and the effect of annealing on junction resistances is important in building large-scale superconducting quantum processors. Here we study the effects of aging of Josephson junctions under different storage conditions from immediately after fabrication up to 2 to 3 months. We find that the aging curve follows a logarithmic curve, with the aging amplitude mainly determined by fabrication conditions and the aging speed determined by storage conditions. Junctions stored at ambient laboratory conditions aged faster compared to junctions stored in a nitrogen atmosphere or vacuum, with the aging speed appreciably changes when the storage condition changed. We also compared the effect of thermal annealing under nitrogen environment with annealing under ambient conditions up to 250C. We find that under nitrogen environment, the resistances decreased at all temperatures tested, while under ambient environment the resistances increased at 200C and decreased at 250C instead. We were unable to decrease the resistance below the initial-time resistance, suggesting a lower limit on the range of resistance tuning.
@@ I Introduction
Josephson junctions are essential elements in superconducting quantum processors [2, 7]. For all types of superconducting qubits, their frequencies depend on the critical current of the junctions, which in turn depends on fabrication conditions including the junction area and oxidation parameters. At room temperature, can be estimated by measuring the junction resistance , which is related to via the Ambegaokar-Baratoff relation [1]. For large-scale quantum processors, precise and accurate junction fabrication is important for qubit frequency assignment to avoid frequency collisions that can limit qubit gate performance [15]. Current state-of-the-art fabrication performance gives about 4% wafer-scale coefficient-of-variation (CV) and 1% chip-scale CV [13].
@@
To improve the frequency accuracy, several junction anneal methods that change the critical current post-fabrication are developed. This include thermal annealing (i.e heating), which has been reported to decrease (increase ) or increase (decrease ) depending on the temperature and the heating environment [22, 11, 12]. However, thermal annealing is limited to chip-scale adjustment only, while junction-scale annealing will be necessary for frequency accuracy improvement. In recent years, laser annealing [8, 25, 10], electron-beam annealing [3], and voltage annealing [19, 24, 9] are developed that allows junction-scale resistance tuning, although they are so far limited to decreasing .
@@
An additional factor that needs to be taken into account is the fact that the junctions “aged”, where the resistance slowly increased over time [11, 18]. As a result, a qubit’s measured frequency can differ significantly from its estimated frequency if its is measured significantly before it is cooled down. Recent reports show that the aging is affected by junction design and fabrication, including junction dimensions and oxidation parameters [14]. Several methods have been shown to inhibit aging, including plasma cleaning [20], additional oxidation steps [4], or using etching instead of lift-off process [23]. However, reducing the aging appears to also reduce the junction response to annealing [14, 4], suggesting a balance between the two is needed in the manufacturing of large scale quantum processors.
@@
In this work, we study the junction aging properties under different storage environments over several months, including ambient atmosphere, nitrogen glove box, and vacuum. We observed that aging speed mainly depend on the storage conditions, with ambient conditions resulted in fastest aging. On the other hand, the aging amplitude appears to be limited mainly by fabrication conditions, similar to recent reports [14]. Changing the storage conditions of the chips resulted in an apparent change in aging speed. We also studied the effect of prior storage on both voltage annealing and thermal annealing of Josephson junctions, with no significant difference observed. The results from voltage annealing experiments suggest this process did not induce an accelerated aging of junctions, but instead changed the internal structure of the junctions. For thermal annealing in nitrogen atmosphere, we only observed decrease in up to 250 C temperatures, consistent with previous reports in low-oxygen environment [22, 12]. For thermal annealing in ambient conditions, we observed increase in for 200 C and decrease in for 250 C. Additionally, the lower limit for resistance decrease appeared to be the initial (immediately after fabrication) value. Understanding junction aging and annealing characteristics allows one to plan the optimal storage and the timing and annealing method for different resistance targets.
@@ II Experimental Details
The test junctions used in this work were fabricated using the Dolan junction fabrication procedure widely used in the fabrication of superconducting qubits [16, 17]. Each test chip consists of 16 junctions, with a typical fabricated size of approximately 250 nm 250 nm.
High-resistivity ( cm) Silicon wafers are diced into individual 1.5 cm 1.5 cm square chips. We spin-coated bilayer e-beam resist, consisting of 500 nm MMA-EL9 and 300 nm PMMA-A4-950K. After drawing the test junction patterns using e-beam lithography, the chips are developed using MIBK:IPA 1:3 solution for 40 s, and then loaded into the e-beam deposition system.
Before deposition, we performed in-situ cleaning and Si removal using a gentle argon ion mill (beam voltage 250 V) [21] at and tilt angles. A small flow of oxygen (Ar flow 9 sccm, O2 flow 0.9 sccm) is included to remove organic residue. Two layers of Al are deposited with 0.5 nm/s rate, with different thicknesses and deposition tilt angles: 1st layer 60 nm and , and 2nd layer 100 nm and . We performed an oxidation step between the deposition of the 2 layers, with 0.5 mbar oxygen pressure and varying duration depending on target resistance. After the second deposition, we performed a final protective oxidation step for 10 mbar and 10 minutes. We performed liftoff using heated acetone (about 50C) for approximately 45 minutes. This liftoff duration was shorter than the typical duration (several hours or overnight) so that we can measure the junction resistance as close to the deposition time as possible. We also did not perform additional bandage layer fabrication [6]. As a result, after the junction deposition the chips were not exposed to any baking step that may result in any significant thermal annealing. The initial resistance measurement of test junctions are typically done within 20 minutes after the end of lift-off (corresponding to about 4 hours after oxidation step), then the chips are stored in their assigned environment.
@@ III.1 Single Environment Chips
The resistances of junctions on Chip 1 (stored in atmosphere) and Chip 2 (stored in glove box), which were simultaneously deposited, were measured over a period of 56 days. Fig. 1(a) shows the individual and averaged resistance of the test junctions during the measurement period, with defined as approximately the time the chips are taken out from the acetone bath at the end of lift-off. Fig. 1(b) show the resistance CV of the resistances over the measurement period. For Chip 1 it grew slowly to about , while for Chip 2 it remained about the same.
@@
where corresponds to the fractional aging amplitude, corresponds to aging speed. We needed to include additional fit parameter that takes into account the initial resistance, and we expect . Table 2 shows the fit parameters of the average resistance for the 3 chips, as well as the average of the fit parameters from individual fits for each junction on each chip. The distribution of the individual junction fit parameters is shown in Fig. 1(c-f). As is located within the logarithmic term of Eq. 1 and the histogram showed an asymmetric distribution in Fig. 1(c), it is reasonable to show the distribution of instead [Fig. 1(d)]. The histograms for parameters and show some overlap for chips 1 and 2, while only show minimal overlap.
@@ IV.1 Alternating Bias Voltage Annealing
| Chip | Junction | (k) | () |
| 1 | Annealed | ||
| Non-Annealed | |||
| 2 | Annealed | ||
| Non-Annealed |
After the aging measurements described in Section III.1, at days after junction deposition, we annealed half of the junctions on Chips 1 and 2 using the ABAA annealing method at room temperature [19, 24], with 30 total square pulses (15 positive, 15 negative), each with 0.9 V amplitude and 1 s duration. Table 3 summarizes the effect of annealing process on the junctions. The non-annealed junctions were not affected by the annealing of nearby junctions. The annealed junctions in both chips increased by an average of approximately 7.5 k, which corresponds to about fractional increase for Chip 1 and about increase for Chip 2.
Afterwards, we measured the post-annealed resistances for another 6 days, as shown in Fig. 3. For the non-annealed junctions on Chip 1, their resistances follow the expected resistance (based on fit parameters in Table 2), while there was a small increase for Chip 2. This is likely due to the fact that during the annealing process, the Chip was out in atmosphere for approximately 90 minutes, 4 to 5 times longer than the typical time it was in the atmosphere during the aging measurements in the previous section, and was under an accelerated aging for a short period. The annealed junctions on both chips follow a similar resistance curve, with initial apparent aging before slowing down after several days, similar to previous reports [24, 14].
@@ V Microscopic picture of resistance aging and annealing
Third, the alternating-bias voltage annealing results (Table 3 and Fig. 3), which show an abrupt increase in for annealed junctions followed by renewed slow drift, can be interpreted within Eq. 3 as a perturbation primarily to the internal reservoir. A voltage pulse of amplitude applied across an oxide of thickness produces an electric field that is large on atomic scales for nanometer-scale oxides. Such fields can drive field-assisted hopping and rearrangement of charged defects (e.g., oxygen vacancies and trapped charge) in amorphous AlOx, effectively “reconfiguring” the dominant conduction paths and increasing the effective barrier parameter in Eq. 2. In this picture, voltage annealing does not simply accelerate the same environment-driven aging mechanism; rather, it changes the internal configuration (shifting and/or ), after which logarithmic relaxation resumes with new initial conditions.
@@ VI Conclusions and Outlook
We also compared the effect of voltage/thermal annealing on junctions with identical fabrication but different storage conditions. The similar resistance change and further logarithmic aging post-voltage annealing suggest that voltage annealing does not “accelerate” the environmental aging process, but instead changed the internal configuration of the junctions. We observed resistance decrease for thermal annealing in N2 environment for all temperatures used (up to 250 C). For thermal annealing in ambient conditions, we observed a slight increase at 200 C and decrease at 250 C, suggesting a competition between deaging and aging (e.g oxygen desorbption/absorption) processes during annealing.
@@ Acknowledgements.
The authors thank Senthil Kumar Karuppannan, Christoph Hufnagel, and Yuanzheng Paul Tan for assistance and valuable discussions. This project is supported by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009) and by the National Research Fund (NRF) Quantum Engineering Program (QEP) W21Qpd0211 Advanced Quantum Processor Platform.References
@@
- [8] (2021) Laser-annealing Josephson junctions for yielding scaled-up superconducting quantum processors. 7 (1), pp. 129. External Links: ISSN 2056-6387, Document, Link Cited by: §I.
@@
- [12] (2024) Aluminum Josephson junction microstructure and electrical properties modified by thermal annealing. Scientific Reportsnpj Quant. Info.Appl. Phys. Lett.Scientific ReportsScientific ReportsAppl.Phys. Lett.Science AdvancesCommunications MaterialsPhys. Rev. Mater.Phys. Rev. BJ. Appl. Phys.Journal of Applied PhysicsApplied Physics LettersApplied Physics ExpressNaturePhys. Rev. Res.Science AdvancesPhys. Rev. Lett.ACS Appl. Mater. InterfacesPhys. Rev. Lett.Rep. Prog. Phys.Phys. Chem. Chem. Phys.The Journal of Chemical PhysicsJ. Phys. Chem. Cphysica status solidi (b)Applied Physics LettersAppl. Phys. Lett.Quant. Sci. Tech.Applied Physics LettersPhys. Rev. Lett.NatureNaturePhys. Rev. Res.Phys. Rev. Lett.Supercond. Sci. Tech. 14 (1), pp. 26066. External Links: ISSN 2045-2322 Cited by: §I, §I, item 2, item 3.

### v2

@@ Abstract
Understanding the aging behavior of Josephson junctions and the effect of annealing on junction resistances is important in building large-scale superconducting quantum processors. Here we study the effects of aging of Al/AlOx/Al Josephson junctions under different storage conditions from immediately after fabrication up to 2 to 3 months. We find that the aging curve follows a logarithmic curve, with the aging amplitude mainly determined by fabrication conditions and the aging speed determined by storage conditions. Junctions stored at ambient laboratory conditions aged faster compared to junctions stored in a nitrogen atmosphere or vacuum, with the aging speed appreciably changes when the storage condition changed. We also compared the effect of thermal annealing under nitrogen environment with annealing under ambient conditions up to 250C. We find that under nitrogen environment, the resistances decreased at all temperatures tested, while under ambient environment the resistances increased at 200C and decreased at 250C instead. We were unable to decrease the resistance below the initial-time resistance, suggesting a lower limit on the range of resistance tuning.
@@ I Introduction
Josephson junctions are essential elements in superconducting quantum processors [2, 7]. For all types of superconducting qubits, their frequencies depend on the critical current of the junctions, which in turn depends on fabrication conditions including the junction area and oxidation parameters. At room temperature, can be estimated by measuring the junction resistance , which is related to via the Ambegaokar-Baratoff relation [1]. For large-scale quantum processors, precise and accurate junction fabrication is important for qubit frequency assignment to avoid frequency collisions that can limit qubit gate performance [8, 15]. Current state-of-the-art fabrication performance gives about 3% wafer-scale coefficient-of-variation (CV) and 1-2% chip-scale CV [13, 23, 28].
@@
To improve the frequency accuracy, several junction anneal methods that change the critical current post-fabrication are developed. These include thermal annealing (i.e heating), which has been reported to decrease (increase ) or increase (decrease ) depending on the temperature and the heating environment [22, 11, 12]. However, thermal annealing is limited to chip-scale adjustment only, while junction-scale annealing will be necessary for frequency accuracy improvement. In recent years, laser annealing [8, 27, 10], electron-beam annealing [3], and voltage annealing [19, 26, 9] are developed that allows junction-scale tuning, although they are so far limited to decreasing .
@@
An additional factor that needs to be taken into account is the fact that the junctions “aged”, where the resistance slowly increased over time [11, 18]. As a result, a qubit’s measured frequency can differ significantly from its expected frequency if its is measured significantly before it is cooled down. Recent reports show that junction aging is affected by junction design and fabrication, including junction dimensions and oxidation parameters [14]. Several methods have been shown to inhibit aging, including plasma cleaning [20], additional oxidation steps [4], or using etching instead of lift-off process [25]. However, reducing the aging appears to also reduce the junction sensitivity to annealing [14, 4], suggesting a balance between the two is needed in the manufacturing of large scale quantum processors.
@@
In this work, we study the Al/AlOx/Al junction aging properties under different storage environments over several months, including ambient atmosphere, nitrogen glove box, and vacuum. We observed that aging speed mainly depend on the storage conditions, with ambient conditions resulted in fastest aging. On the other hand, the aging amplitude appears to be limited mainly by fabrication conditions, similar to recent reports [14]. Changing the storage conditions of the chips results in an apparent change in aging speed, with a resistance decrease ( junction “deaging”) observed when the junctions are moved from ambient atmosphere to glove box. We also studied the effect of prior storage on both voltage annealing and thermal annealing of Josephson junctions. The results from voltage annealing experiments suggest this process did not induce an accelerated aging of junctions, but instead changed the internal structure of the junctions. For thermal annealing in nitrogen atmosphere, we only observed decrease in up to 250 C temperatures, consistent with previous reports in low-oxygen environment [22, 12]. For thermal annealing in ambient conditions, we observed increase in for 200 C and decrease in for 250 C. Additionally, the lower limit for resistance decrease appeared to be the initial (immediately after fabrication) value. Understanding junction aging and annealing characteristics allows one to plan the optimal storage and the timing and annealing method for different resistance targets.
@@ II Experimental Details
The test Al/AlOx/Al junctions used in this work were fabricated using the Dolan junction fabrication procedure widely used in the fabrication of superconducting qubits [16, 17]. Each test chip consists of 16 junctions, with a typical fabricated size of approximately 250 nm 250 nm.
High-resistivity silicon wafers ( cm) are diced into individual 1.5 cm 1.5 cm square chips. We spin-coated bilayer e-beam resist, consisting of 500 nm MMA-EL9 and 300 nm PMMA-A4-950K. After drawing the test junction patterns using e-beam lithography, the chips are developed using MIBK:IPA 1:3 solution for 40 s, and then loaded into the e-beam deposition system.
Before deposition, we performed in-situ cleaning and SiO2 removal using a gentle argon ion mill (beam voltage 250 V) [21] at and tilt angles. A small flow of oxygen (Ar flow 9 sccm, O2 flow 0.9 sccm) is included to remove organic residue. Two layers of Al are deposited with 0.5 nm/s rate, with different thicknesses and deposition tilt angles: 1st layer 60 nm and , and 2nd layer 100 nm and . We performed an oxidation step between the deposition of the 2 layers, with 0.5 mbar oxygen pressure and varying duration depending on target resistance. After the second deposition, we performed a final protective oxidation step for 10 mbar and 10 minutes. We performed liftoff using heated acetone (about 50C) for approximately 45 minutes. This liftoff duration was shorter than the typical duration (several hours or overnight) so that we can measure the junction resistance as close to the deposition time as possible. We also did not perform additional bandage layer fabrication [6]. As a result, after the junction deposition the chips were not exposed to any baking step that may result in any significant thermal annealing. The initial resistance measurement of test junctions are typically done within 20 minutes after the end of lift-off (corresponding to about 4 hours after oxidation step), then the chips are stored in their assigned environment.
@@ III.1 Single Environment Chips
The resistances of junctions on Chip 1 (stored in ambient atmosphere) and Chip 2 (stored in glove box), which were simultaneously deposited, were measured over a period of 56 days. Fig. 1(a) shows the individual and averaged resistance of the test junctions during the measurement period, with defined as approximately the time the chips are taken out from the acetone bath at the end of lift-off. Fig. 1(b) show the resistance CV of the resistances over the measurement period. For Chip 1 it grew slowly to about , while for Chip 2 it remained about the same.
@@
where corresponds to the fractional aging amplitude, corresponds to aging speed. We include an offset parameter inside the logarithm to ensure the value remains finite at , and we expect since the argument of the logarithm should approach unity at early times. Table 2 shows the fit parameters of the average resistance for the 3 chips, as well as the average of the fit parameters from individual fits for each junction on each chip. The distribution of the individual junction fit parameters is shown in Fig. 1(c-f). As is located within the logarithmic term of Eq. 1 and the histogram showed an asymmetric distribution in Fig. 1(c), it is reasonable to show the distribution of instead [Fig. 1(d)]. The histograms for parameters and show some overlap for chips 1 and 2, while only show minimal overlap.
@@ IV.1 Alternating Bias Voltage Annealing
| Chip | Voltage | |||
| Anneal | (k) | (%) | ( s) | |
| 1 | yes | 5.0 | ||
| no | - | |||
| 2 | yes | 1.6 | ||
| no | 1.8 |
After the aging measurements described in Section III.1, at days after junction deposition, we annealed half of the junctions on Chips 1 and 2 using the ABAA annealing method at room temperature [19, 26], with 30 total square pulses (15 positive, 15 negative), each with 0.9 V amplitude and 1 s duration. Table 3 summarizes the effect of annealing process on the junctions. The non-annealed junctions were not affected by the annealing of nearby junctions. The annealed junctions on both chips increased by an average of approximately 7.5 k, which corresponds to about resistance increase for Chip 1 and about increase for Chip 2.
Afterwards, we measured the post-annealed resistances for another 6 days, as shown in Fig. 3. For the non-annealed junctions on Chip 1, their resistances follow the expected resistance (red dashed line, extrapolated to days based on fit parameters in Table 2), while there was a small increase for Chip 2. This is likely due to the fact that during the annealing process, the chip was out in atmosphere for approximately 90 minutes, 4 to 5 times longer than the typical time it was in the atmosphere during the aging measurements in the previous section, and was under an accelerated aging for a short period. The annealed junctions on both chips follow a similar resistance curve, with initial apparent aging before slowing down after several days, similar to previous reports [26, 14]. We fit the resistances of the annealed junctions and the resistances of the non-annealed junctions of Chip 2 to the same logarithmic function of Eq. 1, with defined as the voltage annealing time. Table 3 also shows the characteristic time from the fit, although we expect significant uncertainty due to relatively limited data. That values differ from the post-fabrication time (Table 2) suggests that they do not correspond to aging speed, but possibly relaxation towards a bound curve (see Section III.2), which, for very aged junctions, can be considered almost constant for several days.
@@ V Microscopic picture of resistance aging and annealing
Third, the alternating-bias voltage annealing results (Table 3 and Fig. 3), which show an abrupt increase in for annealed junctions followed by renewed slow drift, can be interpreted within Eq. 3 as a perturbation primarily to the internal reservoir. A voltage pulse of amplitude applied across an oxide of thickness produces an electric field that is large on atomic scales for nanometer-scale oxides. Such fields can drive field-assisted hopping and rearrangement of charged defects (e.g., oxygen vacancies and trapped charge) in amorphous AlOx [24], effectively “reconfiguring” the dominant conduction paths and increasing the effective barrier parameter in Eq. 2. In this picture, voltage annealing does not simply accelerate the same environment-driven aging mechanism; rather, it changes the internal configuration (shifting and/or ), after which the aging resumes with new initial conditions towards a new “bound” curve.
@@ VI Conclusions and Outlook
We also compared the effect of voltage and thermal annealing on junctions with identical fabrication but different storage conditions. The similar change and further aging post-voltage annealing suggest that voltage annealing does not “accelerate” the environmental aging process, but instead changed the internal configuration of the junctions. We observed resistance decrease for thermal annealing in N2 environment for all temperatures used (up to 250 C). For thermal annealing in ambient conditions, we observed a slight increase at 200 C and decrease at 250 C, suggesting a competition between deaging and aging (e.g oxygen desorbption/absorption) processes during annealing.
@@ Acknowledgements.
The authors thank Senthil Kumar Karuppannan, Christoph Hufnagel, and Yuanzheng Paul Tan for their assistance and useful discussions, and D. Scott Holmes for valuable suggestions. This project is supported by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009) and by the National Research Fund (NRF) Quantum Engineering Program (QEP) W21Qpd0211 Advanced Quantum Processor Platform.References
@@
- [8] (2021) Laser-annealing Josephson junctions for yielding scaled-up superconducting quantum processors. 7 (1), pp. 129. External Links: ISSN 2056-6387, Document, Link Cited by: §I, §I.
@@
- [12] (2024) Aluminum Josephson junction microstructure and electrical properties modified by thermal annealing. Scientific Reportsnpj Quant. Info.Appl. Phys. Lett.Scientific ReportsScientific ReportsAppl.Phys. Lett.Science AdvancesCommunications MaterialsPhys. Rev. Mater.Phys. Rev. BJ. Appl. Phys.Journal of Applied PhysicsApplied Physics LettersApplied Physics ExpressNaturePhys. Rev. Res.Science AdvancesPhys. Rev. Lett.ACS Appl. Mater. InterfacesPhys. Rev. Lett.Rep. Prog. Phys.Phys. Chem. Chem. Phys.The Journal of Chemical PhysicsJ. Phys. Chem. Cphysica status solidi (b)Applied Physics LettersAppl. Phys. Lett.Quant. Sci. Tech.Appl. Phys. Lett.Phys. Rev. Lett.NatureNaturePhys. Rev. Res.Phys. Rev. Lett.Supercond. Sci. Tech.Jpn. J. Appl. Phys.Scientific Reports 14 (1), pp. 26066. External Links: ISSN 2045-2322 Cited by: §I, §I, item 2, item 3.
@@
- [23] (2022-12) Uniformity improvement of Josephson-junction resistance by considering sidewall deposition during shadow evaporation for large-scale integration of qubits. 62 (SC), pp. SC1002. Cited by: §I.
- [24] (2025) Simulating alternating bias assisted annealing of amorphous oxide tunnel junctions. External Links: 2512.18420, Link Cited by: §V.
@@
- [28] (2023) Fabrication of Al/AlOx/Al junctions with high uniformity and stability on sapphire substrates. 13 (1), pp. 11874. External Links: ISSN 2045-2322 Cited by: §I.